What mass of melamine, C3N3(NH2)3, will be obtained from 146.0 kg of urea, CO(NH2)2, if the yield of the overall reaction is 73.0 % ?
1
Expert's answer
2017-09-19T12:04:07-0400
6(NH2)2CO = C3H6N6 +3CO2 + 6NH3
M((NH2)CO) = 60g/mol
M (C3H6N6) = 126g/mol
360g (NH2)2CO - 126g C3H6N6
146Kg (NH2)2CO - xKg C3H6N6
x = 146*126/360 = 51.1 Kg C3H6N6 (100% yield)

73% yield = 51.1*0.73 = 37.3 Kg C3H6N6